25 bags of wheat of each weighing 40 kg cost Rs. 2750. Find the cost of 35 bags if it weighs 50kg

Grade:6

2 Answers

Khimraj
6 years ago
There are 25 bags of wheat of each weighing 40kg cost Rs. 2750.
So total weight of wheat is 25*40 = 1000 kg
So cost per kg = 2750/1000 = Rs 2.75
Since total weight of 35 bags each weight of 50kg = 35*50 = 1750 kg
 So cost of these 35 bags are 1750*2.75 = Rs 4812.5
 
Hope it clears. If still  you have doubt you can ask me. If you like answer then please approve it.
Om
5 years ago
Dear student
You can solve this question by simple unitary method (Or)
40 kg will cost you Rs 2750
1kg will cost you (2750/40)=Rs 68.75
35 kg will cost you (35*68.75) =Rs2406.25
 
I hope this answer helps you
With regards 
A fellow student